# GaussianFormer3D: Multi-Modal Gaussian-based Semantic Occupancy Prediction with 3D Deformable Attention

### v1

@@ 3.1 Scene as 3D Gaussian Representation
Gaussian-to-voxel splatting module is designed to only aggregate Gaussians within a neighborhood of the targeted voxel instead of querying all Gaussians in a scene to improve efficiency and reduce unnecessary computation and storage [25]. Thus, Eq. 3 can be furthur approximated by replacing $N_{g}$ with $N_{g}(\mathbf{x})$, where $N_{g}(\mathbf{x})$ is the number of neighboring Gaussians at location $\mathbf{x}$. During training, the Gaussian-based occupancy model is trained in an end-to-end manner, supervised by the ground truth semantic occupancy label $\mathbf{\bar{O}}\in\mathcal{C}^{X\times Y\times Z}$. Both cross entropy loss $L_{ce}$ and the lovasz-softmax loss $L_{lov}$ [2] are used for supervision.
3.2 Voxel-to-Gaussian Initialization
Two sets of 3D Gaussian features are adopted in our model following GaussianFormer [25]. The first set consists of learnable Gaussian physical properties $\mathcal{G}=\{\mathbf{G}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N_{g}}$ introduced in Sec. 3.1, where $d=11+|\mathcal{C}|$, which are also our learning targets. The second set is non-learnable high-dimensional Gaussian features $\mathcal{Q}=\{\mathbf{Q}_{i}\in\mathbb{R}^{m}\}_{i=1}^{N_{g}}$, where $m$ is the feature dimension, serving as queries for the attention mechanism [65, 7, 94] and implicitly encode the spatial and semantic information during the Gaussian update. Previous work [25] randomly initializes the Gaussian physical properties during training, and optimizes these properties iteratively through multiple repetitive refinement modules. This design constrains Gaussians to learn complex 3D geometry information solely from 2D images, which will inevitably encounter inaccurate spatial modeling.
To resolve this issue, we propose a LiDAR-based voxel-to-Gaussian initialization strategy to initialize the mean and opacity of Gaussians with geometry priors from LiDAR data, as indicated in the dashed blue box in Fig. 2. Specifically, we first aggregate the most recent $N_{f}$ LiDAR scans into a combined point cloud $\mathcal{\bar{P}}=\{\mathcal{P}_{i}\}_{i=1}^{N_{f}}$. We voxelize the combined point cloud and compute the feature of each non-empty voxel as the mean position and intensity of all points within it [93]. These LiDAR-based voxel features are then used to initialize the position and opacity of 3D Gaussians:
where $i\in\{1,...,N_{g}\}$ denotes the index of Gaussians to be initialized and $v\in\{1,...,N_{v}\}$ denotes the index of all non-empty voxels; $\mathcal{P}_{v}$ is the set of LiDAR points in $\mathcal{\bar{P}}$ falling into voxel $v$. When $N_{g}<N_{v}$, we randomly choose a subset of non-empty voxels with size of $N_{g}$ to initialize Gaussians, otherwise, we randomly select a subset of Gaussians with size of $N_{v}$ to be initialized with non-empty voxels. After initialization, we apply a 3D sparse convolution module to 3D Gaussians for self-encoding. The information and interactions of these Gaussians are efficiently extracted and aggregated through the sparse convolution operation for updating the Gaussian queries.
3.3 LiDAR-Guided 3D Deformable Attention
Lift, Splat, Shoot (LSS) [54] and 2D attention-based methods [94, 7] are widely adopted for feature lifting which transform multi-view 2D images into a 3D space to obtain lifted features. However, LSS suffers from excessive computational costs, hindering its application to multi-scale feature maps that are important for recognizing objects of various sizes. GaussianFormer [25] utilizes a 2D deformable attention to extract visual information from 2D images. Despite its efficiency, the 2D deformable attention-based method struggles with depth ambiguity. As multiple 3D reference points of different Gaussians can be projected to the same 2D position with similar sampling points in the 2D view, leading to the aggregation of the same 2D features for different 3D Gaussian queries. The underlying reason for this is the lack of accurate depth information during the feature lifting and aggregating process.
A 3D deformable attention operator, namely DFA3D [31], is designed to mitigate the depth ambiguity problem by first expanding 2D feature maps into 3D using estimated depth [12] and then applying an attention mechanism [65] to aggregate features from the expanded 3D feature maps. However, the operator was originally designed for BEV-based 3D object detection, and relies on DepthNet [12] to estimate monocular depth. Inspired by DFA3D [31], we propose a LiDAR-guided 3D deformable attention mechanism for Gaussian-based semantic occupancy prediction, as illustrated in the dashed orange box in Fig. 2. We first form a unified LiDAR-camera 3D feature space $\mathbf{F}^{{\rm 3D}}$ by conducting outer product between the multi-scale depth maps $\mathbf{F}^{\rm d}$, generated from the LiDAR point cloud, and the multi-scale camera feature maps $\mathbf{F}^{\rm c}:\mathbf{F}^{\rm 3D}=\mathbf{F}^{\rm d}\otimes\mathbf{F}^{\rm c}$. For feature sampling, we design a two-stage key point sampling method to aggregate sufficient informative features for updating Gaussian queries. First, we sample a group of 3D reference points $\mathcal{R}_{G}=\{\mathbf{m}_{i}=\mathbf{m}+\Delta\mathbf{m}_{i}|i=1,...,N_{R_{1}}\}$ for each Gaussian $\mathbf{G}$ by shifting its mean $\mathbf{m}$ with learned offsets $\Delta\mathbf{m}$. Then we project these 3D reference points into the fusion feature space $\mathbf{F}^{\rm 3D}$ with extrinsics $\mathcal{T}$ and intrinsics $\mathcal{K}$, where each projected reference point is positioned at $\mathbf{\bar{m}}_{i}=(u_{i},v_{i},d_{i})$. After projection, we further generate learnable sampling offsets $\Delta\mathbf{\bar{m}}_{ij}=(\Delta u_{ij},\Delta v_{ij},\Delta d_{ij})$ for each projected reference point $\mathbf{\bar{m}}_{i}$. The overall sampling points of a given Gaussian $\mathbf{G}$ in the fusion feature space $\mathbf{F}^{\rm 3D}$ can be formulated as:
[…]
@@ 3.3 LiDAR-Guided 3D Deformable Attention
where $N_{R_{1}}$ and $N_{R_{2}}$ denote the number of sampling reference points for each Gaussian and for each projected 3D reference point, respectively. Finally, we update the Gaussian query $\mathbf{Q}$ with the weighted sum of aggregated LiDAR-camera fusion features $\Delta\mathbf{Q}$:
@@
where $\rm DFA(\cdot)$ and $\boldsymbol{\pi}_{c}(\cdot)$ represent the 3D deformable attention operation and the transformation from the Gaussian coordinate frame to $\mathbf{F}_{c}^{{\rm 3D}}$ coordinate frame generated from camera view $c$ respectively. After acquiring sufficient 3D geometric and semantic information through sparse convolution and 3D deformable attention, the Gaussian query $\mathbf{Q}$ is passed to a multi-layer perceptron (MLP), and decoded to refine the Gaussian property $\mathbf{G}$. We iteratively optimize the Gaussian properties with $B$ blocks of sparse convolution, 3D deformable attention, and refinement modules.
@@
Table 1: 3D semantic occupancy prediction results on SurroundOcc [70] validation set. The best is bolded and the second best is underlined. Our method achieves comparable overall performance with the state-of-the-art LiDAR-camera fusion-based methods, while surpasses them on classes of small objects, dynamic vehicles, and large surfaces.
@@
Table 2: 3D semantic occupancy prediction results on Occ3D [64] validation set. * denotes training with camera visibility mask. (xf) denotes the number of history image frames used for temporal fusion. The best is bolded and the second best is underlined.
4 Experiments
4.1 Datasets
@@
RELLIS-3D [27] dataset is a multi-modal off-road driving dataset containing RGB images, LiDAR point clouds, stereo images, GPS and IMU data. WildOcc [83] provides the first off-road occupancy annotation on the RELLIS-3D, which are split into 7399/1249/1399 frames for training, validation and testing respectively. The annotation is in the range of $[-20\rm{m},0\rm{m}]\times[-10\rm{m},10\rm{m}]\times[-2\rm{m},6\rm{m}]$, where each voxel has a resolution of $0.2\mathrm{m}$ and labeled as one of 9 classes (7 semantics, 1 other class and 1 empty). WildOcc [83] is used to evaluate the performance of our model on off-road scenes with a LiDAR-monocular setting.
4.2 Implementation and Evaluation Details
For camera branch, we set the resolution of input images as $900\times 1600$ for nuScenes [4] and $1200\times 1920$ for RELLIS-3D [27]. We utilize the ResNet101-DCN [15] checkpoint pretrained from FCOS3D [67] as the backbone and FPN [40] as the neck. For LiDAR branch, we aggregate and voxelize previous 10 sweeps of point clouds, and obtain the mean features through a voxel feature encoder [93]. The LiDAR depth map is generated and saved before training following [31, 35]. The number of Gaussians is set to 25,600 in our main experiments. We employ these Gaussians to only model the occupied space, and leave the empty space to one fixed large Gaussian to improve efficiency [23]. We train our model with an AdamW [42] optimizer with a weight decay of 0.01. The learning rates are set to $1\times 10^{-4}$ for nuScenes and $3\times 10^{-4}$ for RELLIS-3D, and decay with a cosine annealing schedule. Our model is trained for 24 epochs with a batch size of 8 on nuScenes and 20 epochs with a batch size of 4 on RELLIS-3D on Nvidia A40 GPUs. We use Intersection-over-Union (IoU) and mean Intersection-over-Union (mIoU) for evaluation metrics following MonoScene [5]. See supplementary material for the calculation details of the metrics.
@@ 4.2 Implementation and Evaluation Details
Table 3: 3D semantic occupancy prediction results on WildOcc [83] validation and test sets. We show the percentages of classes in the ground truth occupied voxels for validation and test sets respectively. Results except for GaussianFormer [25] and ours are reported in WildOcc [83]. (xf) denotes the number of history image frames used for temporal fusion. The best is bolded.
@@
BEVFormer [37] C 2D BEV 200$\times$200 310 4.5 30.5 16.8
TPVFormer [22] C 3D TPV 200$\times$(200+16+16) 320 5.1 30.9 17.1
SurroundOcc [70] C 3D Voxel 200$\times$200$\times$16 340 5.9 31.5 20.3
@@
50$\times$50$\times$4 532 7.6 33.3 21.2
M-CONet [68] L+C 3D Voxel 100$\times$100$\times$8 670 7.8 39.2 24.7
Co-Occ [52] L+C 3D Voxel 100$\times$100$\times$8 580 11.8 41.1 27.1
@@
Table 4: Efficiency comparison of different methods and effect of Gaussian number on SurroundOcc [70] validation set. The results of our method are tested on one A40 GPU with batch size one during inference. The best is bolded within each modality.
4.3 Quantitative Results
@@
3D semantic occupancy prediction performance. We report the performance of GaussianFormer3D on SurroundOcc [70], Occ3D [64] and WildOcc [83] in Tab. 1, Tab. 2 and Tab. 3, respectively. For on-road scenarios in Tab. 1 and Tab. 2, our method surpasses GaussianFormer [25] extensively on all classes, leading to overall 13.5 and 8.0 increases on the IoU and mIoU respectively on SurroundOcc [70] and 10.9 increase on the mIoU on Occ3D [64]. Compared to state-of-the-art LiDAR-camera approaches [52, 68, 87], GaussianFormer3D achieves comparable overall performance while showing superior performance in predicting small objects (e.g., motorcycle, pedestrian), dynamic vehicles (e.g., car, construction vehicle, truck) and surrounding surfaces (e.g., manmade, vegetation) which are crucial classes for autonomous driving tasks. This improvement is due to Gaussians’ universal approximating ability to model objects with flexible scales and shapes. For off-road results in Tab. 3, our method with single-frame image input surpasses M-OFFOcc [83] using 4 sequential images by 1.1 in IoU and performs on par in mIoU. Moreover, our method outperforms GaussianFormer [25] by 14.4 in IoU and 6.8 in mIoU on the test set, highlighting LiDAR’s role in understanding the geometry of complex off-road terrains. GaussianFormer3D excels in predicting regions with large surfaces, such as grass, tree, and puddle, while remaining suboptimal for subtle terrain variations like mud. For barrier and rubble, their low occurrence in the test set (0.001% of occupied voxels) poses a challenge due to the lack of sufficient features for reliable prediction. See supplementary material for more analysis.
@@
Evaluation of model efficiency and effect of Gaussian number. We evaluate and compare the latency and memory consumption of our approach with other methods in Tab. 4. Our model achieves multi-modal fusion-based prediction performance while maintaining approximately the same low memory usage as camera-only methods. Compared to Co-Occ [52], our method saves about 50% memory consumption, making it more suitable for running onboard on autonomous vehicles. Additionally, our approach employs only 25,600 Gaussians with 28 channels while Co-Occ [52] requires 80,000 queries with 128 channels to achieve similar performance demonstrating the potential of our method to enable more efficient communication for connected vehicles or multi-robot collaborations. The latency of our method is higher than that of vision-based pipelines, which is mainly due to the computation overhead introduced by 3D deformable attention operations. We also examine the effect of the number of Gaussians on the model performance in Tab. 4. As the number of Gaussians increases, both latency and memory consumption rise, while the IoU and mIoU metrics are steadily improved.
4.4 Ablation Study
To break down the performance improvement brought by two designed modules, we conduct extensive ablation experiments to validate our design choices. The main ablation study is conducted in Tab. 5. We observe that both the proposed voxel-to-Gaussian initialization and the LiDAR-guided 3D deformable attention modules contribute to the superior performance of GaussianFormer3D. The voxel-to-Gaussian initialization significantly improves the model’s ability to detect both small objects (e.g., pedestrian, traffic cone) and large surfaces (e.g., manmade structures, vegetation). This validates the effectiveness of multi-sweep LiDAR scans in providing Gaussians with accurate geometric information of occupied space. We also notice that LiDAR-guided 3D deformable attention mechanism enhances the model’s prediction ability on dynamic vehicles (e.g., bicycle, bus, car, motorcycle, trailer, truck) and near-road surfaces (e.g., drivable surface, flatten area, sidewalk, terrain) where objects detected by LiDAR points are visible to surrounding cameras. In these regions, the LiDAR points and corresponding image pixels are associated in the lifted 3D feature space, enabling the model to retrieve aggregated fusion features of on-road and near-road objects.
@@ 4.4 Ablation Study
Figure 3: Qualitative results on the on-road SurroundOcc [70] validation set. Our multi-modal Gaussian-based occupancy method can capture both semantics information and geometry structure of the surroundings. Best viewed on screen and in color.
@@
Table 5: Ablation study of proposed modules evaluated on SurroundOcc [70] validation set. Voxel-to-Gaussian and LiDAR-Guided 3D Deformable Attention are abbreviated as V2G and DFA respectively.
@@
(a) Ablation study of initialization strategy for V2G. PM-Point denotes probabilistic modeling with point cloud reported in [23].
@@
(d) Ablation study of feature lifting and aggregating methods for DFA. We concatenate LiDAR sparse and dense depth maps with RGB features respectively to conduct 2D deformable attention-based Gaussian update.
Table 6: Ablation study of module design choices on the SurroundOcc [70] validation set. Voxel-to-Gaussian and LiDAR-Guided 3D Deformable Attention are abbreviated as V2G and DFA respectively.
@@
Figure 4: Qualitative results on the off-road WildOcc [83] test set. Our multi-modal Gaussian-based occupancy method can outperform the ground truth (as shown in the first row) and predict classes such as puddle that are vital for off-road autonomous driving (as shown in the second row). Best viewed on screen and in color.
@@
Figure 5: Visualization comparison with GaussianFormer [25] on SurroundOcc [70]. By incorporating LiDAR, our method can obtain Gaussians with more adaptive scales and shapes, resulting in more accurate semantic predictions and delicate geometry details.
Figure 6: Multi-resolution occupancy prediction using same 3D Gaussians. Leveraging the continuity of Gaussians, our method enables multi-resolution prediction, yielding more accurate and smoother occupancy at higher resolution in certain regions.

### v2

@@ III-C LiDAR-Guided 3D Deformable Attention
where $N_{R_{1}}$ and $N_{R_{2}}$ denote the number of sampling points for each Gaussian and for each projected 3D reference point. Finally, the Gaussian query $\mathbf{Q}$ is updated with the weighted sum of aggregated LiDAR-camera fusion features $\Delta\mathbf{Q}$:
@@
where $\rm DFA(\cdot)$ and $\boldsymbol{\pi}_{c}(\cdot)$ represent the 3D deformable attention and the transformation from the Gaussian frame to $\mathbf{F}_{c}^{{\rm 3D}}$ frame generated from camera view $c$, respectively. After acquiring sufficient geometric and semantic information through sparse convolution and 3D deformable attention, the Gaussian query $\mathbf{Q}$ is passed to a multi-layer perceptron, and decoded to refine the Gaussian property $\mathbf{G}$. We iteratively optimize the properties with 4 blocks of sparse convolution, 3D deformable attention, and refinement modules.
TABLE I: 3D semantic occupancy prediction results on nuScenes-SurroundOcc [38] validation set.
@@
TABLE II: 3D semantic occupancy prediction results on nuScenes-Occ3D [32] validation set. * denotes training with camera visibility mask. (xf) denotes the number of history image frames used for temporal fusion.
@@
IV Experiments
IV-A Datasets
@@
RELLIS-3D [12] dataset is a multi-modal off-road driving dataset collected by a Clearpath Warthog robot containing RGB images, LiDAR point clouds, stereo images, GPS and IMU data. WildOcc [44] provides the first off-road occupancy annotation on the RELLIS-3D, which are split into 7399/1249/1399 frames for training, validation and testing respectively. The annotation is in the range of $20\rm{m},0\rm{m}]\times[-10\rm{m},10\rm{m}]\times[-2\rm{m},6\rm{m}]$, where each voxel has a resolution of $0.2\mathrm{m}$ and labeled as one of 9 classes (7 semantics, 1 other class and 1 empty). WildOcc [44] is used to evaluate the performance of our model in complex off-road environments and with a monocular-LiDAR sensor configuration.
IV-B Implementation and Evaluation Details
For camera branch, we set the resolution of input images as $900\times 1600$ for nuScenes [2] and $1200\times 1920$ for RELLIS-3D [12]. We utilize the ResNet101-DCN [7] checkpoint pretrained from FCOS3D [35] as the backbone and FPN [21] as the neck. For LiDAR branch, we aggregate and voxelize previous 10 sweeps of point clouds, and obtain the mean features through a voxel feature encoder [51]. The LiDAR depth map is generated and saved before training following [17, 18]. The number of Gaussians is set to 25,600 in our main experiments. We employ these Gaussians to only model the occupied space, and leave the empty space to one fixed large Gaussian to improve efficiency [9]. We train our model with an AdamW optimizer with a weight decay of 0.01. The learning rates are set to $1\times 10^{-4}$ for nuScenes and $3\times 10^{-4}$ for RELLIS-3D, and decay with a cosine annealing schedule. Our model is trained for 24 epochs with a batch size of 8 on nuScenes and 20 epochs with a batch size of 4 on RELLIS-3D on Nvidia A40 GPUs. We use Intersection-over-Union (IoU) and mean Intersection-over-Union (mIoU) for evaluation metrics following MonoScene [3].
TABLE III: 3D semantic occupancy prediction results on RELLIS3D-WildOcc [44] dataset.
@@ IV-B Implementation and Evaluation Details
TABLE IV: Efficiency comparison and ablation on number of Gaussians. Tested on one A40 GPU with one batch during inference.
@@
BEVFormer [19] C 2D BEV $\times200$ 310 4.5 30.5 16.8
[…]
SurroundOcc [38] C 3D Voxel $\times200$$\times$16 340 5.9 31.5 20.3
@@
$\times50$$\times$4 532 7.6 33.3 21.2
M-CONet [36] L+C 3D Voxel $\times100$$\times$8 670 7.8 39.2 24.7
Co-Occ [26] L+C 3D Voxel $\times100$$\times$8 580 11.8 41.1 27.1
@@
IV-C Quantitative Results
@@
3D semantic occupancy prediction performance. We report the performance of GaussianFormer3D on SurroundOcc [38], Occ3D [32] and WildOcc [44] in Tab. I, Tab. II and Tab. III, respectively. For on-road scenarios in Tab. I and Tab. II, our method surpasses GaussianFormer [11] extensively on all classes, leading to overall 13.5 and 8.0 increases on the IoU and mIoU respectively on SurroundOcc [38] and 10.9 increase on the mIoU on Occ3D [32]. Compared to state-of-the-art LiDAR-camera approaches [26, 36, 45], ours achieves best overall performance while showing superior performance in predicting small objects (e.g., motorcycle, pedestrian), dynamic vehicles (e.g., car, construction vehicle, truck) and surrounding surfaces (e.g., manmade, vegetation) which are crucial classes for autonomous driving tasks. This improvement is due to Gaussians’ universal approximating ability to model objects with flexible scales and shapes. For off-road results in Tab. III, our method with single-frame image input surpasses M-OFFOcc [44] using 4 sequential images by 1.1 in IoU and performs on par in mIoU. Moreover, our method outperforms GaussianFormer [11] by 14.4 in IoU and 6.8 in mIoU on the test set, highlighting LiDAR’s role in understanding the geometry of complex off-road terrains. Our method excels in predicting regions with large surfaces, such as grass, tree, and puddle, while remaining suboptimal for subtle terrain variations like mud. For barrier and rubble, their low occurrence in the test set (0.001% of occupied voxels) poses a challenge due to the lack of sufficient features for reliable prediction. We further evaluate the model performance under different weather conditions in Tab. V. Ours shows a significant performance improvement over the baseline under extreme climate (rainy) and low lighting condition (night).
@@
Evaluation of model efficiency. We evaluate and compare the latency and average memory consumption of ours with other methods during testing in Tab. IV. Ours achieves multi-modal fusion-based prediction performance while maintaining approximately the same low memory usage as camera-only methods. Compared to Co-Occ [26], ours saves about 50% average memory consumption, making it more suitable for running onboard on autonomous vehicles. In addition, our approach employs only 25,600 Gaussians with 28 channels while Co-Occ [26] requires 80,000 queries with 128 channels to achieve similar performance, demonstrating the potential of our method to enable more efficient communication for connected vehicles or multi-robot collaborations. The latency of our method is higher than that of camera-only pipelines, which is mainly due to the computation overhead introduced by 3D deformable attention. Some LiDAR-camera methods [45, 44] are not compared due to lack of open-source code. We also examine the effect of the number of Gaussians on the model performance in Tab. IV. As the number of Gaussians increases, both latency and memory consumption rise, while the IoU and mIoU metrics are steadily improved.
@@
Figure 4: Qualitative results on the on-road nuScenes-SurroundOcc [38] validation set. Our multi-modal Gaussian-based occupancy method can capture both semantics information and geometry structure of the surroundings. Color legend is given in Tab. I.
TABLE V: Performance on nuScenes-SurroundOcc [38] validation set under different weather and lighting conditions.
[…]
Method Modality [Uncaptioned image] Sunny [Uncaptioned image] Rainy [Uncaptioned image] Day [Uncaptioned image] Night [Uncaptioned image] Sunny [Uncaptioned image] Rainy [Uncaptioned image] Day [Uncaptioned image] Night
GaussianFormer [11] C 29.6 27.5 30.3 19.5 18.9 18.0 19.2 9.3
TABLE VI: Ablation study of proposed modules evaluated on nuScenes-SurroundOcc [38] validation set. Voxel-to-Gaussian and LiDAR-Guided 3D Deformable Attention are abbreviated as V2G and DFA respectively.
@@
TABLE VII: Ablation study of module design choices on the nuScenes-SurroundOcc [38] validation set.
(a) Ablation study of Gaussian initialization strategies. PM-
Point denotes probabilistic modeling with point cloud in [9].
@@
(d) Ablation study of feature lifting and aggregating methods for DFA. We concatenate LiDAR sparse and dense depth maps with RGB features respectively to conduct 2D deformable attention.
@@
Figure 5: Qualitative results on the off-road RELLIS3D-WildOcc [44] test set. Our method can outperform the GT (first row) at some regions and predict classes such as puddle that are vital for off-road autonomous driving (second row). Color legend is given in Tab. III.
@@
Figure 6: Visualization comparison with GaussianFormer [11] on nuScenes-SurroundOcc [38]. By incorporating LiDAR, our method can obtain Gaussians with more adaptive scales and shapes, resulting in more accurate semantic predictions and delicate geometry details.
Figure 7: Multi-resolution occupancy prediction of 3D Gaussians.
IV-D Ablation Study
We conduct extensive ablation experiments to validate our design choices. The main ablation study is conducted in Tab. VI. We observe that both the proposed voxel-to-Gaussian initialization and the LiDAR-guided 3D deformable attention modules contribute to the superior performance of our method. The voxel-to-Gaussian initialization significantly improves the model’s ability to detect both small objects (e.g., pedestrian, traffic cone) and large surfaces (e.g., manmade, vegetation). This validates the effectiveness of multi-sweep LiDAR scans in providing Gaussians with accurate geometric information of occupied space. We also notice that LiDAR-guided 3D deformable attention mechanism enhances the model’s prediction ability on dynamic vehicles (e.g., bicycle, bus, car, motorcycle, trailer, truck) and near-road surfaces (e.g., drivable surface, flatten area, sidewalk, terrain) where objects detected by LiDAR points are visible to surrounding cameras. In these regions, the LiDAR points and corresponding image pixels are associated in the lifted 3D feature space, enabling the model to retrieve aggregated fusion features of on-road and near-road objects.